Fermi National Laboratory
 
Operation Summary for November 23, 2011 - November 28, 2011
 

NuMI reported receiving 8.31E18 protons on target for the period from 11/21/11 to 11/28/11.

Wednesday November 23

The day shift began with Operations monitoring the Main Injector (MI), Linac, and Booster, and with the MiniBooNE and NuMI taking beam.

During the day shift, the Meson FTBF T-978 experiment began taking beam at 11 AM. Muon and MI personnel began conducting studies around noon. At 12:41 PM, operators prepared Booster for an access to replace a power amplifier for a RF station (BRF1). NuMI also accessed during this downtime. Booster resumed sending beam at 3:40 PM.

During the evening shift, operations disabled beam to T-978 at 7:11 PM.

Thursday November 24

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, operators rebooted a NuMI BPM VME (NBPM6N) crate at 1:15 AM. A MI RF station (MIRF15) tripped off for the first of three times in four hours; it reset.

During the day shift, operators reported at 3:17 PM that MIRF15 had tripped off multiple times during the shift. Operators reported short downtimes due to LRF1 and LRF2 trips.

During the evening shift, operators rebooted a NuMI BPM (NBPM8N) at 5:39 PM. LRF5 tripped off for the first of seven times in about two hours; it reset. Operators reported after the last trip that all three switch tubes were showing arcing indications. MIRF15 tripped off at 11:50 PM; it reset.

Friday November 25

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 4:43 AM, MIRF15 tripped off for the first of three times in about 20 minutes; it reset. At 6:50 AM, LRF5 tripped off due to a blown fuse; operators changed the fuse, but it blew again. Operators contacted Linac personnel. Operators loaded an energy conservation timeline, at 7:30 AM, into the Time Line Generator (TLG).

During the day shift, Linac personnel began troubleshooting LRF5; during the shift they replaced the SCR control cards, the modulator’s switch tubes, an overload relay, and resistors. At the end of the shift, they were slowly running the station up, but more problems occurred.

During the evening shift, LRF5 repairs continued through out the shift.

Saturday November 26

The midnight shift began with no beam due to LRF5 repairs.

During the midnight shift, Linac personnel brought LRF5 online at 12:40 AM. Operations established beam to NuMI and MiniBooNE around 12:55 AM. Operations reported at 4:22 AM that due to an I- Source current drop, they had raised the Source’s arc supply voltage by one volt. Operations established beam to a new Meson FTBF experiment, T-1012 at 7 AM: they will only be running during this weekend. T-978 will resume taking beam on Monday.

During the day shift, Operations monitored the accelerator complex.

During the evening shift, Operations disabled beam to T-1012 at 6:58 PM. Operators completed a Booster injection tune at 7:09 PM. At 7:23 PM, operators reported repeated trips by BRF15 that caused the NuMI beam permit to also trip off; operators bypassed the station.

Sunday November 27

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, Operations established beam to T-1012 at 7:27 AM.

During the day shift, at 1:20 PM, LRF1 tripped off for the first of two times in about 40 minutes; the station would not reset after the second trip. Operators called in Linac personnel and loaded an energy conservation timeline into the TLG at 2:14 PM.

During the evening shift, Linac personnel reported at 4:15 PM that the LRF1’s driver tube needed replacing. T-1012 decided to end their experimental run at 7:34 PM. LRF1 came back online and Operations established beam to MiniBooNE and NuMI by 7:56 PM. LRF1 tripped off at 8:28 PM; it reset.

Monday November 28

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 1:12 AM, the NuMI beam permit tripped off due an absorber room water skid indication; Operations contacted a NuMI person. He reported that there was nothing wrong with the water skid, but thought it was a controls issue with the PLC. Operators talked with a Controls person who had operators run some tests that pointed to a problem with the PLC and not controls. Operations disabled the NuMI beam, increased the beam intensity for MiniBooNE, and called in an Engineer to look at the PLC. At 5:48 AM, a MI Recycler kicker (KPS3A) tripped off on a low Fluorinert indication. Operations turned on the Muon Rings, Transport, PreTarget, and PreVault areas at 7:46 AM.

Operations Coordinator (OpsCo)

The OpsCo approved of the plans by Booster and MI personnel to conduct studies and to continue the NuMI PLC replacement.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/28/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory